OpenStudy (anonymous):

Which mixed number is equivalent to the decimal plotted at point G on the number line?

OpenStudy (anonymous):

Which mixed number is equivalent to the decimal plotted at point G on the number line? Number line from 4.4 to 6, tic marks every 0.1 unit, label even decimals from 4.4 through 6, dot at 5.4 with label G above the dot. A. B. C. D.

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

5.4 = (27/5) = 5 + (2/5)